Determine the charge in each capacitor and the voltage across each capacitor.

Given: V3= 5V

C1=1uF

C2= 1uF

C3= 1uF

C4= 1uF

C5= 1uF


1
Expert's answer
2021-05-18T17:30:18-0400

Equivalent Capacitance, "C_{eq}=\bigg(\dfrac{1}{C_1}+\dfrac{1}{C_2}+\dfrac{1}{C_3}+\dfrac{1}{C_4}+\dfrac{1}{C_5}\bigg)^{-1}"

"C_{eq}=\dfrac{1}{5}=0.20\space\mu F"

Voltage across "c_3,\space V_3=5\space V"

Therefore charge stored in "c_3,\space q_3=c_3\times V_3=5\times10^{-6}\space C"

Since, the capacitors are connected in series therefore charge stored in each capacitor is same and so is the voltage as all capacitors have same capacitance i.e "q_1=q_2=q_3=q_4=q_5=5\times10^{-6}\space C"

and

"V_1=V_2=V_3=V_4=V_5=5\space V"


Total voltage, "V=25\space V"